# A Lie algebraic pattern behind logarithmic CFTs

### v2

@@ 2.2. Simplicity theorem
We will omit the detail of the proof by the same reason above. Let us consider the later half. Using the Leibniz rule (2.1) repeatedly, $W_{P+\lambda_{\bullet}}$ is generated by $|\lambda_{\bullet}\rangle$ as $W_{P}$-module. In the same manner as Lemma 2.7, $W_{P+\lambda_{\bullet}}$ is simple as $W_{P}$-module (by the same discussion for $W_{Q}$ and $W_{Q+\lambda_{\bullet}}$, the first half of the Lemma is already proved at this point). On the other hand, $W_{Q+\lambda}$ is generetad by some element in $L_{\lambda^{\bullet}}\otimes\mathbb{C}^{\times}|\lambda\rangle$ as $W_{Q}$-module. Since $W_{Q+\lambda}\subseteq W_{P+\lambda_{\bullet}}$ and the $W_{P+\lambda_{\bullet}}$ is simple as $W_{P}$-module, by [DM97, Corollary 4.2], any element of $W_{Q+\lambda}$ has the form $\sum_{n\in\mathbb{Z}}X^{n}_{(n)}|\lambda\rangle$ for some $\{X^{n}\}_{n\in\mathbb{Z}}\subset W_{P}$. By weight consideration, we can show that $\{X^{n}\}_{n\in\mathbb{Z}}\subset W_{Q}$, and thus $W_{Q+\lambda}$ is simple as $W_{Q}$-module. ∎
Let us explain how these results will be used to the simplicity theorems of (multiplet) $W$-(super)algebras. We consider the case (1) or (2). Then, by Lemma 2.7, $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module. To extend the simplicity of $\mathcal{W}_{-\beta}$ to $\mathcal{W}_{-\beta+\lambda}$ such that $\lambda\in\Lambda$ satisfies (strong), we consider using Lemma 2.6 to $\lambda_{0}=0$ and $\lambda_{1}=\lambda$. Here, $\mathbb{M}(\mu)$ and $\mathbb{L}(\mu)$ represent the Verma module and the irreducible module over $\mathcal{W}_{0}$ with highest weight $\mu$, respectively. As we show later, sometimes $\mathbb{L}(-\beta+\lambda_{i})$ ($i=0,1$) has a ”Kazhdan-Lusztig-type” character formula in Lemma 2.6(3). Since $\mathcal{W}_{-\beta+\lambda_{0}}$ is irreducible, by Lemma 2.6, so is $\mathcal{W}_{-\beta+\lambda_{1}}$. In particular, $\mathcal{W}_{-\beta+\lambda}$ are cyclic. Furthermore, if they are CFT-type, by Lemma 2.8, $W_{Q+\lambda_{\bullet}}$ or $W_{Q+\lambda}$ is also simple as $W_{Q}$-module.
In [Sug21, Sug22] and [CNS24], the cases where $\mathcal{W}_{0}$ is the principal $W$-algebra $\mathcal{W}_{k}(\mathfrak{g})$ for simply-laced $\mathfrak{g}$ and affine $\mathfrak{sl}_{2}$ $V_{k}(\mathfrak{sl}_{2})$ were considered, respectively, and the corresponding simplicity theorems are proved in the same manner explained above. In Section 3 and Section 4 of the present paper, we consider the cases where $\mathcal{W}_{0}$ is the principal $W$-(supr)algebra $\mathcal{W}_{k}(\mathfrak{g})$ for non-simply-laced $\mathfrak{g}$ and $\mathfrak{g}=\mathfrak{osp}(1|2r)$, respectively. In the former case, (2) holds and the simplicity theorem will be proved for any $\lambda\in\Lambda$ satisfying (strong), but in the later case, it will be proved only for $\lambda^{\bullet}=0$ case (i.e., assume only (1)) because of some technical difficulty.
@@ 2.3.1. Fermionic VOSA
$Y(\phi(-\tfrac{1}{2})\mathbb{1},z)=\phi(z)=\sum_{n\in\frac{1}{2}+\mathbb{Z}}\phi(n)z^{-n-\frac{1}{2}},\ \ \omega^{(s)}=\tfrac{1}{2}\phi(-\tfrac{3}{2})\phi(-\tfrac{1}{2})\mathbb{1},$ (2.4)
@@
$\operatorname{ch}_{q}F=\tfrac{\eta(q)^{2}}{\eta(q^{2})\eta(q^{\frac{1}{2}})},\ \ \operatorname{sch}_{q}F=\tfrac{\eta(q^{\frac{1}{2}})}{\eta(q)},\ \ \operatorname{ch}\iota^{\ast}F=2\tfrac{\eta(q^{2})}{\eta(q)}\ \ \text{($\iota$ is the parity isomorphism)},$) ,
@@ 2.3.2. Lattice VOSAs
where $h\in\hat{\mathfrak{h}}_{L}$ acts as $(\mu,h)$ on $\mathbb{C}_{\mu}$. Then there exists a simple V(S)A structure (see [Kac98]) on
@@
where $e^{\mu}(e^{\lambda})=e^{\mu+\lambda}$ for $\lambda\in\mathbb{C}[{L}]$, and $\epsilon(\cdot,\cdot)$ is the $2$-cocycle satisfying [Kac98, (5.4.14)]. The parity of the element $v\otimes e^{\mu}$ is $|\mu|^{2}\mod{2}$. One can define a (shifted) conformal vector
$\omega_{\mathrm{sh}}=\omega_{\mathrm{st}}+\gamma_{(-2)}\mathbf{1},\ \ \omega_{\mathrm{st}}=\tfrac{1}{2}\sum_{i=1}^{r}v_{i(-1)}v_{i}^{\ast},\ \ \gamma\in\mathfrak{h}_{L}$ (2.5)
with central charge is $r-12(\gamma,\gamma)$, where $\{v^{i}\}_{i=1}^{\ell}$ and $\{v_{i}^{\ast}\}_{i=1}^{r}$ are bases of $L$ and $L^{\ast}$, respectively, such that $(v_{i},v_{j}^{\ast})=\delta_{i,j}$. The dual lattice $L^{\ast}$ of $L$ is defined by $L^{\ast}:=\{\beta\in\mathbb{Q}\otimes L\;|\;(\alpha,\beta)\in\mathbb{Z}\,\,\text{for all}\,\,\alpha\in L\}$. Then the lattice VO(S)A has finitely many irreducible modules parametrized by $L^{\ast}/L$ [LiL04]; these modules are
@@
$\Delta_{\mu}=\tfrac{1}{2}|\mu-\gamma|^{2}+\tfrac{c-r}{24}=\tfrac{1}{2}|\mu|^{2}-(\mu,\gamma).$ (2.6)
According to [FHL93, Proposition 5.3.2], the contragredient dual $V_{L+\lambda}^{\ast}$ of $V_{L+\lambda}$ is $V_{L+\lambda^{\prime}}$, for some $\lambda^{\prime}\in L^{\ast}/L$.
@@ 2.4. Construction of shift systems
$h_{i}=\lceil-\tfrac{1}{x}\alpha_{i(0)}^{\vee}\rceil.$ (2.7)
@@
$V_{\lambda}=V_{\sqrt{p}(Q+\lambda)},\ \ \omega=\omega_{\mathrm{sh}},\ \ h_{i}=\lceil-\tfrac{1}{\sqrt{p}}\alpha_{i(0)}^{\vee}\rceil,\ \ f_{i}=e^{\sqrt{p}\alpha_{i}}_{(0)}\ \ (1\leq i\leq r),$ (2.8)
$V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}\otimes F,\ \ \omega=\omega_{\mathrm{sh}}+\omega^{(s)},\ \ h_{i}=\lceil-\tfrac{1}{\sqrt{p}}\alpha_{i(0)}^{\vee}\rceil,\ \ f_{i}=\begin{cases}e^{\sqrt{p}\alpha_{i}}_{(0)}&(1\leq i\leq r-1),\\ (e^{\sqrt{p}\alpha_{r}}\otimes\phi)_{(0)}&(i=r),\end{cases}$ (2.9)
@@ 2.4.1. The case (1)
By (2.6), the operators $f_{i}$ ($i\in\Pi$) preserve the conformal weight iff
$\omega=$ $\tfrac{1}{2}\sum_{i\in\Pi}\alpha_{i(-1)}\alpha_{i}^{\ast}+\sqrt{p}(\rho-\tfrac{1}{p}\rho^{\vee})_{(-2)}\mathbf{1}$ (2.10)
and thus we consider this conformal vector. On the other hand, under the choice (2.10), another solution of $\Delta_{e^{x\alpha_{i}}_{(0)}}=0$ is $x=-\tfrac{2}{\sqrt{p}(\alpha_{i},\alpha_{i})}$, namely the case $x\alpha_{i}=-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}$. By [TK86], if $(p\lambda+\rho^{\vee},\alpha_{i})\equiv s\ (\operatorname{mod}\ p)$ for $11\leq s\leq p-$ 1, then we have the non-zero (conformal weight-preserving) screening operators
[…]
where $[\Gamma]$ is the cycle defined in [NT11, Section 2.3]. To use (2.11) as short screening operators of a shift system, we have to regard $\lambda-\tfrac{s}{p}\alpha_{i}^{\vee}$ as $\sigma_{i}\ast\lambda$ for some $W$-action $\ast$. If we choose $x=\tfrac{1}{p}\rho^{\vee}$ in Example 1.8, then the representative $\Lambda$ of $\tfrac{1}{p}Q^{\ast}/Q$ is given by
$\Lambda=\{-\lambda^{\bullet}+\lambda_{\bullet}\ |\ \lambda^{\bullet}\in P_{\mathrm{min}},\ \lambda_{\bullet}=\sum_{i\in\Pi}\tfrac{\lambda_{i}-1}{p}\alpha_{i}^{\ast},\ 1\leq\lambda_{i}\leq r_{i}^{\vee}m\},\ r_{i}^{\vee}=\begin{cases}r^{\vee}&(\text{$\alpha_{i}$ is long}),\\ 1&(\text{$\alpha_{i}$ is short}),\end{cases}$ (2.12)
@@
$\sigma\ast\mu=\sigma(\mu+\tfrac{1}{p}\rho^{\vee})-\tfrac{1}{p}\rho^{\vee}.$ (2.13)
@@ 2.4.2. The case (2)
By (2.6) and $\Delta_{\phi}=\tfrac{1}{2}$, the operators $f_{i}$ ($i\in\Pi$) preserve the conformal weight iff
$\omega=\omega^{(s)}+\tfrac{1}{2}\sum_{i\in\Pi}\alpha_{i(-1)}\alpha_{i}^{\ast}+\sqrt{p}(1-\tfrac{1}{p})\rho_{(-2)}\mathbf{1}$ (2.14)
and thus we consider this conformal vector (note that the $\omega_{\mathrm{sh}}$-part is different from (2.10)). On the other hand, under the choice (2.14), another solution of $\Delta_{e^{x\alpha_{i}}_{(0)}}=0$ ($i\not=r$) and $\Delta_{(e^{x\alpha_{r}}\otimes\phi)_{(0)}}=0$ is $x=-\tfrac{1}{\sqrt{p}}$. In the same manner as above, if $(p\lambda+\rho,\alpha_{i}^{\vee})\equiv s\ (\operatorname{mod}\ p)$ for $11\leq s\leq p-$ 1, then we have the non-zero (conformal weight-preserving) screening operators (see also [IK03])
[…]
@@
[…]
@@
$\sigma\ast\mu=\sigma(\mu+\tfrac{1}{p}\rho)-\tfrac{1}{p}\rho.$ (2.17)
@@
The following lemma is proved in the same manner as [Sug21, Appendix], so we will omit the proof.
@@
The operators $\{f_{i},h_{i}\}_{i\in\Pi}$ give rise to an integrable $\mathfrak{b}$-action on $V_{\lambda}$.
@@
The proof is based on the evaluation of conformal weights. Let us check the Serre relation $(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=0$ ($[h_{i,\lambda},f_{j}]=-c_{ij}f_{j}$ is clear). Set $f_{i}=x_{(0)}$ and $f_{j}=y_{(0)}$. By (2.1), we have $(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=(f_{i}^{1-c_{ij}}y)_{(0)}$ and $f_{i}^{1-c_{ij}}y\in V_{\lambda}^{h=-((1-c_{ij})\alpha_{i}+\alpha_{j})}$. Since $\Delta_{f_{i}}=\Delta_{f_{j}}=0$, we have $\Delta_{f_{i}^{1-c_{ij}}y}=\Delta_{y}=1$. On the other hand, by (2.6) and (2.10), the minimal conformal weight of $V_{\lambda}^{h=-((1-c_{ij})\alpha_{i}+\alpha_{j})}$ is given by
@@
It suffices to check that $[f_{i},Q_{j,\lambda}]=0$. First, we consider the case where $s=1$. By calculation, we have $e^{\sqrt{p}\alpha_{i}}_{(r)}e^{-\frac{1}{\sqrt{p}}\alpha_{j}}=0$ for $r\geq 0$, $i\neq j$. Thus, we have $[f_{i},{Q}_{j,\lambda}]=0$ for $i\neq j$. When $i\neq n$, we have
$e^{\sqrt{p}\alpha_{i}}_{(0)}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}=\sqrt{p}(\alpha_{i})_{(-1)}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{i}}\mathbb{1}=\tfrac{p}{p-1}(L_{-1}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{i}})\mathbb{1}.$ (2.18)
@@
$e^{\sqrt{p}\alpha_{n}}_{(-1)}e^{-\frac{1}{\sqrt{p}}\alpha_{n}}=\tfrac{p}{p-1}(L_{-1}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{n}})\mathbb{1}.$ (2.19)
@@
$[f_{i},e^{-\frac{1}{\sqrt{p}}\alpha_{j}}(z)]=0,\quad\text{for}\;\;i\neq j,$ (2.20)
$[f_{i},e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z)]=\tfrac{p}{p-1}\tfrac{d}{dz}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z),\quad\text{for}\;\;i\neq n,$ (2.21)
$[f_{n},(e^{-\frac{1}{\sqrt{p}}\alpha_{n}}\otimes\phi(-\tfrac{1}{2}))(z)]=\tfrac{p}{p-1}\tfrac{d}{dz}e^{-\frac{1}{\sqrt{p}}\alpha_{n}}(z).$ (2.22)
Note $(L_{-1}u)_{(0)}=0$ in every generalized vertex operator algebra. we conclude that $[f_{i},{Q}_{j,\lambda}]=0$ for $1\leq i\leq n$. Now let us consider the case $s\geq 2$. Using (2.20), we have
@@
$0\rightarrow\ker Q_{i,\lambda}\rightarrow V_{\lambda}\rightarrow\ker Q_{i,\sigma\ast\lambda}(\sigma_{i}\uparrow\lambda)\rightarrow 0$
@@
Let $M(0)^{i,\perp}$ and $M(0)^{i}$ be the rank $r-1$ and rank $1$ Heisenberg VOAs generated by $\{\varpi_{j}|j\not=i\}$ and $\alpha_{i}$, respectively. Then we have the orthogonal decomposition
@@
=:absent:=:= : $M(0)^{i,\perp}\otimes(\bigoplus_{{\alpha}\in Q/\sim}V_{\alpha+\lambda}^{i})$
of $V_{\lambda}$, where $\alpha\sim\beta$ iff $\beta\in\alpha+\mathbb{Z}\alpha_{i}$ and the term $(\otimes F)$ is omitted in the case (1). Since $f_{i}$ and $Q_{i,\lambda}$ commutes with the $M(0)^{i,\perp}$-action, we have the orthogonal decompositions
@@
  • In the cases where $\alpha_{i}$ is long, the assertion follows from [AM08, Theorem 1.1,1.2] for $p=r^{\vee}m$.
@@
Similarly, the last assertion follows from [AM08, Theorem 1.3] and [AM09, Theorem 6.3], respectively. ∎

### v3

@@ 2.2. Simplicity theorem
We will omit the detail of the proof by the same reason above. Let us consider the later half. Using the Leibniz rule (2.1) repeatedly, $W_{P+\lambda_{\bullet}}$ is generated by $|\lambda_{\bullet}\rangle$ as $W_{P}$-module. In the same manner as Lemma 2.7, $W_{P+\lambda_{\bullet}}$ is simple as $W_{P}$-module (by the same discussion for $W_{Q}$ and $W_{Q+\lambda_{\bullet}}$, the first half of the Lemma is already proved at this point). On the other hand, $W_{Q+\lambda}$ is generetad by some element in $L_{\lambda^{\bullet}}\otimes\mathbb{C}^{\times}|\lambda\rangle$ as $W_{Q}$-module. Since $W_{Q+\lambda}\subseteq W_{P+\lambda_{\bullet}}$ and the $W_{P+\lambda_{\bullet}}$ is simple as $W_{P}$-module, by [19, Corollary 4.2], any element of $W_{Q+\lambda}$ has the form $\sum_{n\in\mathbb{Z}}X^{n}_{(n)}|\lambda\rangle$ for some $\{X^{n}\}_{n\in\mathbb{Z}}\subset W_{P}$. By weight consideration, we can show that $\{X^{n}\}_{n\in\mathbb{Z}}\subset W_{Q}$, and thus $W_{Q+\lambda}$ is simple as $W_{Q}$-module. ∎
Let us explain how these results will be used to the simplicity theorems of (multiplet) W-(super)algebras. We consider the case (1) or (2). Then, by Lemma 2.7, $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module. To extend the simplicity of $\mathcal{W}_{-\beta}$ to $\mathcal{W}_{-\beta+\lambda}$ such that $\lambda\in\Lambda$ satisfies (strong), we consider using Lemma 2.6 to $\lambda_{0}=0$ and $\lambda_{1}=\lambda$. Here, $\mathbb{M}(\mu)$ and $\mathbb{L}(\mu)$ represent the Verma module and the irreducible module over $\mathcal{W}_{0}$ with highest weight $\mu$, respectively. As we show later, sometimes $\mathbb{L}(-\beta+\lambda_{i})$ ($i=0,1$) has a ”Kazhdan–Lusztig-type” character formula in Lemma 2.6(3). Since $\mathcal{W}_{-\beta+\lambda_{0}}$ is irreducible, by Lemma 2.6, so is $\mathcal{W}_{-\beta+\lambda_{1}}$. In particular, $\mathcal{W}_{-\beta+\lambda}$ are cyclic. Furthermore, if they are CFT-type, by Lemma 2.8, $W_{Q+\lambda_{\bullet}}$ or $W_{Q+\lambda}$ is also simple as $W_{Q}$-module.
In [45, 46] and [14], the cases where $\mathcal{W}_{0}$ is the principal W-algebra $\mathcal{W}_{k}(\mathfrak{g})$ for simply-laced $\mathfrak{g}$ and the affine VOA $V_{k}(\mathfrak{sl}_{2})$ were considered, respectively, and the corresponding simplicity theorems are proved in the same manner explained above. In Section 3 and Section 4 of the present paper, we consider the cases where $\mathcal{W}_{0}$ is the principal W-(supr)algebra $\mathcal{W}_{k}(\mathfrak{g})$ for non-simply-laced $\mathfrak{g}$ and $\mathfrak{g}=\mathfrak{osp}(1|2r)$, respectively. In the former case, (2) holds and the simplicity theorem will be proved for any $\lambda\in\Lambda$ satisfying (strong), but in the later case, it will be proved only for $\lambda^{\bullet}=0$ case (i.e., assume only (1)) because of some technical difficulty.
@@ 2.3.1. Fermionic VOSA
(2.4) $Y(\phi(-\tfrac{1}{2})\mathbb{1},z)=\phi(z)=\sum_{n\in\frac{1}{2}+\mathbb{Z}}\phi(n)z^{-n-\frac{1}{2}},\ \ \omega^{(s)}=\tfrac{1}{2}\phi(-\tfrac{3}{2})\phi(-\tfrac{1}{2})\mathbb{1},$
@@
$\operatorname{ch}_{q}F=\tfrac{\eta(q)^{2}}{\eta(q^{2})\eta(q^{\frac{1}{2}})},\ \ \operatorname{sch}_{q}F=\tfrac{\eta(q^{\frac{1}{2}})}{\eta(q)},\ \ \operatorname{ch}\iota^{\ast}F=2\tfrac{\eta(q^{2})}{\eta(q)}\ \ \text{($\iota$ is the parity isomorphism)},$
@@ 2.3.2. Lattice VOSAs
where $h\in\hat{\mathfrak{h}}_{L}$ acts as $(\mu,h)$ on $\mathbb{C}_{\mu}$. Then there exists a simple V(S)A structure (see [33]) on
@@
where $e^{\mu}(e^{\lambda})=e^{\mu+\lambda}$ for $\lambda\in\mathbb{C}[{L}]$, and $\epsilon(\cdot,\cdot)$ is the $2$-cocycle satisfying [33, (5.4.14)]. The parity of the element $v\otimes e^{\mu}$ is $|\mu|^{2}\mod{2}$. One can define a (shifted) conformal vector
(2.5) $\omega_{\mathrm{sh}}=\omega_{\mathrm{st}}+\gamma_{(-2)}\mathbf{1},\ \ \omega_{\mathrm{st}}=\tfrac{1}{2}\sum_{i=1}^{r}v_{i(-1)}v_{i}^{\ast},\ \ \gamma\in\mathfrak{h}_{L}$
with central charge is $r-12(\gamma,\gamma)$, where $\{v^{i}\}_{i=1}^{\ell}$ and $\{v_{i}^{\ast}\}_{i=1}^{r}$ are bases of $L$ and $L^{\ast}$, respectively, such that $(v_{i},v_{j}^{\ast})=\delta_{i,j}$. The dual lattice $L^{\ast}$ of $L$ is defined by $L^{\ast}:=\{\beta\in\mathbb{Q}\otimes L\;|\;(\alpha,\beta)\in\mathbb{Z}\,\,\text{for all}\,\,\alpha\in L\}$. Then the lattice VO(S)A has finitely many irreducible modules parametrized by $L^{\ast}/L$ [41]; these modules are
@@
(2.6) $\Delta_{\mu}=\tfrac{1}{2}|\mu-\gamma|^{2}+\tfrac{c-r}{24}=\tfrac{1}{2}|\mu|^{2}-(\mu,\gamma).$
According to [25, Proposition 5.3.2], the contragredient dual $V_{L+\lambda}^{\ast}$ of $V_{L+\lambda}$ is $V_{L+\lambda^{\prime}}$, for some $\lambda^{\prime}\in L^{\ast}/L$.
@@ 2.4. Construction of shift systems
(2.7) $h_{i}=\lceil-\tfrac{1}{x}\alpha_{i(0)}^{\vee}\rceil.$
@@
(2.8) $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)},\ \ \omega=\omega_{\mathrm{sh}},\ \ h_{i}=\lceil-\tfrac{1}{\sqrt{p}}\alpha_{i(0)}^{\vee}\rceil,\ \ f_{i}=e^{\sqrt{p}\alpha_{i}}_{(0)}\ \ (1\leq i\leq r),$
(2.9) $V_{\lambda}=V_{\sqrt{p}(Q+\lambda)}\otimes F,\ \ \omega=\omega_{\mathrm{sh}}+\omega^{(s)},\ \ h_{i}=\lceil-\tfrac{1}{\sqrt{p}}\alpha_{i(0)}^{\vee}\rceil,\ \ f_{i}=\begin{cases}e^{\sqrt{p}\alpha_{i}}_{(0)}&(1\leq i\leq r-1),\\ (e^{\sqrt{p}\alpha_{r}}\otimes\phi)_{(0)}&(i=r),\end{cases}$
@@ 2.4.1. The case (1)
By (2.6), the operators $f_{i}$ ($i\in I$) preserve the conformal weight iff
(2.10) $\omega=$ $\tfrac{1}{2}\sum_{i\in\Pi}\alpha_{i(-1)}\alpha_{i}^{\ast}+\sqrt{p}(\rho-\tfrac{1}{p}\rho^{\vee})_{(-2)}\mathbf{1}$
and thus we consider this conformal vector. On the other hand, under the choice (2.10), another solution of $\Delta_{e^{x\alpha_{i}}_{(0)}}=0$ is $x=-\tfrac{2}{\sqrt{p}(\alpha_{i},\alpha_{i})}$, namely the case $x\alpha_{i}=-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}$. By [49], if $(p\lambda+\rho^{\vee},\alpha_{i})\equiv s\ (\operatorname{mod}\ p)$ for $1\leq s\leq p-1$, then we have the non-zero (conformal weight-preserving) screening operators
(2.11) $Q_{i,\lambda}=\int_{[\Gamma]}e^{-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}}(z_{1})\cdots e^{-\tfrac{1}{\sqrt{p}}\alpha_{i}^{\vee}}(z_{s})dz_{1}...d_{z_{s}}\colon V_{\lambda}\rightarrow V_{\lambda-\tfrac{s}{p}\alpha_{i}^{\vee}},$
where $\Gamma]$ is the cycle defined in [44, Section 2.3]. To use (2.11) as short screening operators of a shift system, we have to regard $\lambda-\tfrac{s}{p}\alpha_{i}^{\vee}$ as $\sigma_{i}\ast\lambda$ for some $W$-action $\ast$. If we choose $x=\tfrac{1}{p}\rho^{\vee}$ in Example 1.8, then the representative $\Lambda$ of $\tfrac{1}{p}Q^{\ast}/Q$ is given by
(2.12) $\Lambda=\{-\lambda^{\bullet}+\lambda_{\bullet}\ |\ \lambda^{\bullet}\in P_{\mathrm{min}},\ \lambda_{\bullet}=\sum_{i\in\Pi}\tfrac{\lambda_{i}-1}{p}\alpha_{i}^{\ast},\ 1\leq\lambda_{i}\leq r_{i}^{\vee}m\},\ r_{i}^{\vee}=\begin{cases}r^{\vee}&(\text{$\alpha_{i}$ is long}),\\ 1&(\text{$\alpha_{i}$ is short}),\end{cases}$
@@
(2.13) $\sigma\ast\mu=\sigma(\mu+\tfrac{1}{p}\rho^{\vee})-\tfrac{1}{p}\rho^{\vee}.$
@@ 2.4.2. The case (2)
By (2.6) and $\Delta_{\phi}=\tfrac{1}{2}$, the operators $f_{i}$ ($i\in I$) preserve the conformal weight iff
(2.14) $\omega=\omega^{(s)}+\tfrac{1}{2}\sum_{i\in I}\alpha_{i(-1)}\alpha_{i}^{\ast}+\sqrt{p}(1-\tfrac{1}{p})\rho_{(-2)}\mathbf{1}$
and thus we consider this conformal vector (note that the $\omega_{\mathrm{sh}}$-part is different from (2.10)). On the other hand, under the choice (2.14), another solution of $\Delta_{e^{x\alpha_{i}}_{(0)}}=0$ ($i\not=r$) and $\Delta_{(e^{x\alpha_{r}}\otimes\phi)_{(0)}}=0$ is $x=-\tfrac{1}{\sqrt{p}}$. In the same manner as above, if $(p\lambda+\rho,\alpha_{i}^{\vee})\equiv s\ (\operatorname{mod}\ p)$ for $1\leq s\leq p-1$, then we have the non-zero (conformal weight-preserving) screening operators (see also [31])
(2.15) $Q_{i,\lambda}=\begin{cases}\int_{[\Gamma]}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z_{1})\cdots e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z_{s})dz_{1}...d_{z_{s}}&(i\not=r),\\ \int_{[\Gamma]}(e^{-\frac{1}{\sqrt{p}}\alpha_{r}}\otimes\phi)(z_{1})\cdots(e^{-\frac{1}{\sqrt{p}}\alpha_{r}}\otimes\phi)(z_{s})dz&(i=r),\end{cases}\colon V_{\lambda}\rightarrow V_{\lambda-\frac{s}{p}\alpha_{i}}.$
@@
(2.16) $\Lambda=\{-\lambda^{\bullet}+\lambda_{\bullet}\ |\ \lambda^{\bullet}\in P_{\mathrm{min}},\ \lambda_{\bullet}=\sum_{i\in I}\tfrac{\lambda_{i}-1}{p}\varpi_{i},\ 1\leq\lambda_{i}\leq p,\ \ \lambda_{r}+(\lambda^{\bullet},\alpha_{r}^{\vee})\equiv 1\ (\operatorname{mod}\ 2)\},$
@@
(2.17) $\sigma\ast\mu=\sigma(\mu+\tfrac{1}{p}\rho)-\tfrac{1}{p}\rho.$
@@
The following lemma is proved in the same manner as [45, Appendix], so we will omit the proof.
@@
The operators $\{f_{i},h_{i}\}_{i\in I}$ give rise to an integrable $\mathfrak{b}$-action on $V_{\lambda}$.
@@
The proof is based on the evaluation of conformal weights. Let us check the Serre relation
$(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=0$
($h_{i,\lambda},f_{j}]=-c_{ij}f_{j}$ is clear). Set $f_{i}=x_{(0)}$ and $f_{j}=y_{(0)}$. By (2.1), we have $(\operatorname{ad}f_{i})^{1-c_{ij}}f_{j}=(f_{i}^{1-c_{ij}}y)_{(0)}$ and $f_{i}^{1-c_{ij}}y\in V_{\lambda}^{h=-((1-c_{ij})\alpha_{i}+\alpha_{j})}$. Since $\Delta_{f_{i}}=\Delta_{f_{j}}=0$, we have $\Delta_{f_{i}^{1-c_{ij}}y}=\Delta_{y}=1$. On the other hand, by (2.6) and (2.10), the minimal conformal weight of $V_{\lambda}^{h=-((1-c_{ij})\alpha_{i}+\alpha_{j})}$ is given by
@@
It suffices to check that $f_{i},Q_{j,\lambda}]=0$. First, we consider the case where $s=1$. By calculation, we have $e^{\sqrt{p}\alpha_{i}}_{(r)}e^{-\frac{1}{\sqrt{p}}\alpha_{j}}=0$ for $r\geq 0$, $i\neq j$. Thus, we have $f_{i},{Q}_{j,\lambda}]=0$ for $i\neq j$. When $i\neq n$, we have
(2.18) $e^{\sqrt{p}\alpha_{i}}_{(0)}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}=\sqrt{p}(\alpha_{i})_{(-1)}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{i}}\mathbb{1}=\tfrac{p}{p-1}(L_{-1}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{i}})\mathbb{1}.$
@@
(2.19) $e^{\sqrt{p}\alpha_{n}}_{(-1)}e^{-\frac{1}{\sqrt{p}}\alpha_{n}}=\tfrac{p}{p-1}(L_{-1}e^{(\sqrt{p}-\frac{1}{\sqrt{p}})\alpha_{n}})\mathbb{1}.$
@@
(2.20) $[f_{i},e^{-\frac{1}{\sqrt{p}}\alpha_{j}}(z)]=0,\quad\text{for}\;\;i\neq j,$
(2.21) $[f_{i},e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z)]=\tfrac{p}{p-1}\tfrac{d}{dz}e^{-\frac{1}{\sqrt{p}}\alpha_{i}}(z),\quad\text{for}\;\;i\neq n,$
(2.22) $[f_{n},(e^{-\frac{1}{\sqrt{p}}\alpha_{n}}\otimes\phi(-\tfrac{1}{2}))(z)]=\tfrac{p}{p-1}\tfrac{d}{dz}e^{-\frac{1}{\sqrt{p}}\alpha_{n}}(z).$
Note $(L_{-1}u)_{(0)}=0$ in every generalized vertex operator algebra. we conclude that $f_{i},{Q}_{j,\lambda}]=0$ for $1\leq i\leq n$. Now let us consider the case $s\geq 2$. Using (2.20), we have
@@
$0\rightarrow\ker Q_{i,\lambda}\rightarrow V_{\lambda}\rightarrow\ker Q_{i,\sigma\ast\lambda}(\sigma_{i}\uparrow\lambda)\rightarrow 0.$
@@
Let $M(0)^{i,\perp}$ and $M(0)^{i}$ be the rank r1r-1 and rank $1$ Heisenberg VOAs generated by $\{\varpi_{j}|j\not=i\}$ and $\alpha_{i}$, respectively. Then we have the orthogonal decomposition
@@
$=:$ $M(0)^{i,\perp}\otimes(\bigoplus_{{\alpha}\in Q/\sim}V_{\alpha+\lambda}^{i})$
of $V_{\lambda}$, where $\alpha\sim\beta$ iff $\beta\in\alpha+\mathbb{Z}\alpha_{i}$ and the term $(\otimes F)$ is omitted in the case 1. Since $f_{i}$ and $Q_{i,\lambda}$ commutes with the $M(0)^{i,\perp}$-action, we have the orthogonal decompositions
@@
  • In the cases where $\alpha_{i}$ is long, the assertion follows from [7, Theorem 1.1,1.2] for $p=r^{\vee}m$.
@@
Similarly, the last assertion follows from [7, Theorem 1.3] and [9, Theorem 6.3], respectively. ∎